Circular No. 6581

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 E. P. Liang, A. Crider, D. Lin, and I. A. Smith, Rice University; and M.

Kusunose, Kwansei-Gakuin University, report: “The x-ray photon fluxes of GRB 970228 and the fading counterpart SAX J0501.7+1146 together follow a t^{−1} decay law (where t represents time) that is a robust prediction of the saturated Comptonization model of γ-ray bursters (Liang et al. 1997, Ap.J. Lett., in press). This model also predicts that the fluxes in other bands from optical to radio, if not absorbed or reprocessed by intervening matter, should follow a similar decay profile. For example, scaling from the x-ray flux, V -band searches of the revised error box (IAUC 6578) deeper than mag 24 + 2.5 log (t/8 hr) may still catch the afterglow.”

R. M. Wagner, Ohio State University, writes: “Examination of R-band CCD images obtained by M. W. Buie with the Lowell Observatory’s 1.1-m and 1.8-m telescopes on Mar. 7.23 ad 8.13 UT confirms the presence of a faint object at R ∼ 21 (Margon et al., IAUC 6577), consistent with the VLA position of a radio source (Frail et al., IAUC 6576) that lies in the BeppoSax Narrow Field x-ray error box (Costa et al., IAUC 6576), and possibly associated with GRB 970228 (IAUC 6572; Palmer et al., IAUC 6577). A PPM solution for the mean position of the optical object yields α = 5^{h}01^{m}41^{s}.70, δ = +11^{o}47′04′′.1 (equinox J2000.0, ± 0′′.2 in both coordinates). We note that the object appears nonstellar on these images and either consists of a condensation with structure encompassing the radio position or it is the close superposition of two distinct sources, one of which is optically fainter and coincident with the radio source. Deep CCD imaging in excellent seeing is required to discriminate between the two cases.”

Wagner and C. B. Foltz, Multiple Mirror Telescope (MMT) Observatory; and P. Hewett, Institute of Astronomy, University of Cambridge, report: “We obtained a noisy spectrum (range 400–970 nm, resolution 2.2 nm) of the object at the position described above with the MMT on Mar. 10.11 UT. We tentatively identify the Ca II H and K break at 592.5 nm, the G-band at 644.7 nm, and the MgH band at 779.8 nm in the reduced spectrum as being consistent with a galaxy of redshift 0.50 ± 0.01. No strong emission lines characteristic of an AGN or QSO were seen. We estimate that V − R = +1.7 from the spectrum. If the rapidly decaying x-ray source observed by the BeppoSax LECS with the flux given on IAUC 6576 is associated with this galaxy, then its luminosity was ∼ 4 × 10^{45} erg/s (assuming H_{0} = 50 km s^{−1} Mpc^{−1}), comparable to the persistent x-ray luminosities of some AGNs and QSOs. Additional spectroscopic observations are required to confirm the redshift of the galaxy.”

1997 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green